An aqueous solution of HF is added to an aqueous solution of K(OH) and a reaction occurs. The

concentration of H + ions is 2.68x10 -1 M. Calculate the concentration of the (OH) - ions. Is the solution

more acidic or basic?



1
Expert's answer
2022-06-10T16:08:04-0400

Solution:

[H+] = 2.68×10−1 M

We can convert between [H+] and pH using the following equations:

pH = −log[H+]

pH = −log(2.68×10−1) = 0.57

pH = 0.57

Any solution that has a pH of less than 7 is considered acidic


For any aqueous solution at 25∘C:

[H+] × [OH−] = Kw = 1.0×10−14

Therefore,

[OH−] = Kw / [H+]

[OH−] = (1.0×10−14) / (2.68×10−1) = 3.73×10−14

[OH−] = 3.73×10−14 M


Answer:

The concentration of hydroxide ions (OH−) is 3.73×10−14 M

The solution is acidic